You slide a 325N trunk up a 20.0° inclined plane with a constant velocity by exerting a force of 211N up, parallel to the inclined plane.

What is the component of the trunk’s weight parallel to the plane?

What is the sum of your applied force, friction and the parallel component of the trunk’s weight?  Why?

What is the size and direction of the frictional force?

What is the coefficient of friction?


1
Expert's answer
2021-05-12T14:59:06-0400

1)


"W_1=325\sin{20}=111\ N"

2)


"F_{net}=0\ N"

3)


"F_f=211-111=100\ N"

Against the motion.


4)


"\mu=\frac{100}{325\cos{20}}=0.327"